# Prioritized Information Bottleneck Theoretic Framework with Distributed Online Learning for Edge Video Analytics

### v2

@@ IV-B Variational Approximation Method
Combining the joint entropy equation (20) with inequality (21), we get:
@@ IV-E Gate Mechanism Based on Distributed Online Learning
The gate mechanism based on distributed online learning is designed to solve the combinatorial multi-armed bandit (CMAB) problem (11) formulated in Sec. III-B. In this subsection, we first introduce the intuitive ideas of gate mechanism. Then, we provide the details and explanation of the pseudocode. Finally, the evaluation of regret performance and communication cost for distributed execution is analyzed mathematically.
@@ IV-E2 Regret Analysis
The regret analysis reflects the efficiency and adaptability of the algorithm in optimizing network performance. A lower regret bound indicates that the algorithm performs close to the optimal strategy, ensuring high inference accuracy and efficient resource utilization despite the dynamic environment and varying network conditions. In the context of a multi-camera sensing network, the choices between different cameras and edge servers are interdependent. Furthermore, the connections between different cameras and edge servers exhibit heterogeneity, with each super arm having different distribution parameters. Therefore, we consider a CMAB problem with a non-identically distributed (non-i.i.d) assumption and derive its regret upper bound in the following part.
The regret $R_{L}(T)$ is an important metric to evaluate the performance of online learning algorithm. The regret over a time horizon $T$ is defined as the difference between the maximum expected reward obtainable by an optimal strategy and the expected reward obtained by the algorithm:
@@
$\begin{split}R(T)\leq O\Bigg{(}&\left(\sqrt{2\sum_{k=1}^{\mathcal{K}^{\text{arm}}}\sigma_{r_{k}}^{2}}+2\mathcal{K}^{\text{arm}}\sqrt{\frac{2}{a_{N}}}\right)\mathcal{K}^{\text{arm}}\sqrt{T\ln T}\\ &+\frac{2\mathcal{K}^{\text{arm}}}{3}\ln T-2\mathcal{K}^{\text{arm}}\sqrt{\frac{2\ln T}{a_{N}}}\Bigg{)},\end{split}$ (28)
where $\mathcal{K}^{\text{arm}}=\sum_{k=K_{\min}}^{K_{\max}}{C_{K}^{k}S^{k+1}}$ represents the maximum number of arms in the optimal super arm, $\sigma_{r_{k}}$ is the standard deviation of the reward for arm $k$, and $a_{N}$ is an upper bound on the linear growth rate of the number of times UCB algorithm arms are selected over time [^6] It indicates that $N_{k}(t)$ follows a linear growth trend, i.e., $N_{k}(t)\approx a_{N}t$.. The non-i.i.d nature of the rewards represents the heterogeneity in the network where different arms can have different reward distributions due to varying network conditions, processing capabilities, and data qualities.
@@
Since $\hat{\mu}_{k,s}(t)$ is an unbiased estimate of $\mu_{k,s}$, its expected value is zero. Thus, the regret is mainly determined by the term $\sqrt{\frac{2\ln t}{N_{k,s}(t)}}$. We approximate the cumulative sum of this term by using an integral, given that $N_{k,s}(t)$ is assumed to grow linearly with time, i.e., $N_{k,s}(t)\approx a_{N}t$, where $a_{N}$ is a constant. Under this assumption, we have:
@@
where $\mathcal{K}^{\text{arm}}$ is the number of arms in the optimal super arm, and $\sigma_{r}^{2}$ is the maximum variance among all arms. Therefore, when $T$ is sufficiently large, the overall regret bound can be expressed as $R(T)\leq O\left(\left(\sqrt{2\sigma_{r}^{2}}+2\sqrt{\frac{2}{a_{N}}}\right)\mathcal{K}^{\text{arm}}\sqrt{T\ln T}+\frac{2\mathcal{K}^{\text{arm}}}{3}\ln T-2\mathcal{K}^{\text{arm}}\sqrt{\frac{2\ln T}{a_{N}}}\right).$ Thus, the cumulative regret $R(T)$ for the distributed UCB algorithm is bounded by the sum of the regret from all camera-edge server pairs, ensuring that the regret grows sub-linearly with respect to $T$. $\blacksquare$
@@ V-A Simulation Setup
  • TOCOM-TEM[16]: A task-oriented communication framework utilizing a temporal entropy model for edge video analytics. It employs the deterministic Information Bottleneck principle to extract and transmit compact, task-relevant features, integrating spatial-temporal data on the server for improved inference accuracy.
@@
To evaluate the impact of communication bottlenecks and delayed cameras on perception accuracy, we present in Fig. 8(a)8(c) the relationships between communication constraints and the perception accuracy. Nevertheless, the benefit of collaborative perception is accompanied by excessive communication overhead. The communication bottleneck refers to network capacity constraints that prevent real-time data transmission, causing frame latency. This issue is prevalent in UDP-based wireless streaming systems, where high throughput often results in out-of-order or delayed frames due to varying channel quality and jitter. Moreover, different coding schemes cause varying delays in dynamic channel conditions, misaligning data fusion due to channel quality and jitter. Therefore, in order to evaluate how latency differences affect perception accuracy, we set communication bottleneck constraints. In our experiments, we use MODA (Multiple Object Detection Accuracy) and MODP (Multiple Object Detection Precision) to assess coding efficiency and robustness.
@@
Our priority-based mechanism selects the camera with the most targets within RoI for the highest transmission priority. Collaboration priority is thus determined by the target count in each camera’s perception area. As shown in Table II, just 0.82 KB of perception data achieves a MODA accuracy of 55.17% and MODP of 69.72%, highlighting significant redundancy among edge cameras. Adding more cameras initially improves perception significantly but offers diminishing returns as communication costs to the edge server increase. In Fig. 11, we show the relationship between communication bottleneck and total latency for different algorithms. By leveraging a priority information bottleneck framework and the UCB algorithm to reduce redundancy, our PIB, despite slightly higher encoding latency than the traditional video codecs, can achieve much lower transmission latency due to its efficient compression. Under a 25 KB/s bottleneck, our PIB reduces latency by 25.1% over TOCOM-TEM and 74.8% over JPEG. At 50 KB/s, our PIB outperforms TOCOM-TEM by 18.3% and JPEG by 65.1%, respectively. At 100 KB/s, PIB achieves 10.8% and 46.1% lower latency than TOCOM-TEM and JPEG, respectively. The encoding latency results of our PIB in different edge devices are presented in Table III. It can be observed that the feature map generation phase dominates the overall encoding latency, while the entropy coding phase contributes a negligible amount of time. Furthermore, edge devices with higher computating capacity exhibit significantly lower encoding latency.
@@ VI Conclusion
In this paper, we have proposed the Prioritized Information Bottleneck (PIB) framework as a robust solution for collaborative edge video analytics. Our contributions are two-fold. First, we have developed a prioritized inference mechanism to intelligently determine the importance of different camera’ FOVs, effectively addressing the constraints imposed by channel capacity and data redundancy. Second, the PIB framework showcases its effectiveness by notably decreasing communication overhead and improving tracking accuracy without requiring video reconstruction at the edge server. Extensive real-world experiments show that: PIB not only surpasses the performance of conventional methods like TOCOM-TEM, JPEG, H.264, H.265, and AV1 with a marked improvement of up to 17.8% in MODA but also achieves a considerable reduction in communication costs by 82.65%, while retaining low latency and high-quality multi-view sensory data processing under less favorable channel conditions.

### v3

@@ IV-B Variational Approximation Method
Combining the joint entropy equation (20) with inequality (21), we obtain:
@@ IV-E Gate Mechanism Based on Distributed Online Learning
The gate mechanism based on distributed online learning is designed to solve the CMAB problem (11) formulated in Sec. III-B. In this subsection, we first introduce the intuitive ideas of gate mechanism. Then, we provide the details and explanation of the pseudocode. Finally, the evaluation of regret performance and communication cost for distributed execution is analyzed mathematically.
@@ IV-E2 Regret Analysis
The regret analysis reflects the efficiency and adaptability of the algorithm in optimizing network performance. A lower regret bound indicates that the algorithm performs close to the optimal strategy, ensuring high inference accuracy and efficient resource utilization despite the dynamic environment and varying network conditions. In the context of a multi-camera sensing network, the choices between different cameras and edge servers are interdependent. Furthermore, the connections between different cameras and edge servers exhibit heterogeneity, with each super arm having different distribution parameters. Therefore, we consider a CMAB problem with a non-identically distributed (non-i.i.d.) assumption and derive its regret upper bound in the following section.
The regret $R_{L}(T)$ is an important metric to evaluate the performance of the online learning algorithm. The regret over a time horizon $T$ is defined as the difference between the maximum expected reward obtainable by an optimal strategy and the expected reward obtained by the algorithm:
@@
$\begin{split}R(T)\leq O\Bigg{(}&\left(\sqrt{2\sum_{k=1}^{\mathcal{K}^{\text{arm}}}\sigma_{r_{k}}^{2}}+2\mathcal{K}^{\text{arm}}\sqrt{\frac{2}{a_{N}}}\right)\mathcal{K}^{\text{arm}}\sqrt{T\ln T}\\ &+\frac{2\mathcal{K}^{\text{arm}}}{3}\ln T\Bigg{)},\end{split}$ (28)
where $\mathcal{K}^{\text{arm}}=\sum_{k=K_{\min}}^{K_{\max}}{C_{K}^{k}S^{k+1}}$ represents the maximum number of arms in the optimal super arm, $\sigma_{r_{k}}$ is the standard deviation of the reward for arm $k$, and $a_{N}$ is an upper bound on the linear growth rate of the number of times UCB algorithm arms are selected over time [^6] It indicates that $N_{k}(t)$ follows a linear growth trend, i.e., $N_{k}(t)\approx a_{N}t$.. The non-i.i.d. nature of the rewards represents the heterogeneity in the network where different arms can have different reward distributions due to varying network conditions, processing capabilities, and data qualities.
@@
Since $\hat{\mu}_{k,s}(t)$ is an unbiased estimate of $\mu_{k,s}$, the expected value of $\hat{\mu}_{k,s}(t)-\mu_{k,s}$ is zero. Thus, the regret is mainly determined by the term $\sqrt{\frac{2\ln t}{N_{k,s}(t)}}$. We approximate the cumulative sum of this term by using an integral, given that $N_{k,s}(t)$ is assumed to grow linearly with time, i.e., $N_{k,s}(t)\approx a_{N}t$, where $a_{N}$ is a constant. Under this assumption, we have:
@@
where $\mathcal{K}^{\text{arm}}$ is the number of arms in the optimal super arm, and $\sigma_{r}^{2}$ is the maximum variance among all arms. Since $O$ notation represents an asymptotic upper bound, the negative term $-2\mathcal{K}^{\text{arm}}\sqrt{\frac{2\ln T}{a_{N}}}$ is omitted to maintain the validity of the bound. Therefore, when $T$ is sufficiently large, the overall regret bound can be expressed as $R(T)\leq O\left(\left(\sqrt{2\sigma_{r}^{2}}+2\sqrt{\frac{2}{a_{N}}}\right)\mathcal{K}^{\text{arm}}\sqrt{T\ln T}+\frac{2\mathcal{K}^{\text{arm}}}{3}\ln T\right).$ Thus, the cumulative regret $R(T)$ for the distributed UCB algorithm is bounded by the sum of the regret from all camera-edge server pairs, ensuring that the regret grows sub-linearly with respect to $T$. $\blacksquare$
@@ V-A Simulation Setup
  • TOCOM-TEM[16]: A task-oriented communication framework utilizing a temporal entropy model for edge video analytics. It employs the deterministic IB principle to extract and transmit compact, task-relevant features, integrating spatial-temporal data on the server for improved inference accuracy.
@@
To evaluate the impact of communication bottlenecks and delayed cameras on perception accuracy, we present in Fig. 8(a)8(c) the relationships between communication constraints and the perception accuracy. Nevertheless, the benefit of collaborative perception is accompanied by excessive communication overhead. The communication bottleneck refers to network capacity constraints that prevent real-time data transmission, causing frame latency. This issue is prevalent in UDP-based wireless streaming systems, where high throughput often results in out-of-order or delayed frames due to varying channel quality and jitter. Moreover, different coding schemes cause varying delays under dynamic channel conditions, misaligning data fusion due to channel quality and jitter. Therefore, in order to evaluate how latency differences affect perception accuracy, we set communication bottleneck constraints. In our experiments, we use MODA (Multiple Object Detection Accuracy) and MODP (Multiple Object Detection Precision) to assess coding efficiency and robustness.
@@
Our priority-based mechanism selects the camera with the most targets within the RoI for the highest transmission priority. Collaboration priority is thus determined by the target count in each camera’s perception area. As shown in Table II, just 0.82 KB of perception data achieves a MODA accuracy of 55.17% and MODP of 69.72%, highlighting significant redundancy among edge cameras. Adding more cameras initially improves perception significantly but offers diminishing returns as communication costs to the edge server increase. In Fig. 11, we show the relationship between communication bottleneck and total latency for different algorithms. By leveraging a priority information bottleneck framework and the UCB algorithm to reduce redundancy, our PIB, despite slightly higher encoding latency than the traditional video codecs, can achieve much lower transmission latency due to its efficient compression. Under a 25 KB/s bottleneck, our PIB reduces latency by 25.1% over TOCOM-TEM and 74.8% over JPEG. At 50 KB/s, our PIB outperforms TOCOM-TEM by 18.3% and JPEG by 65.1%, respectively. At 100 KB/s, PIB achieves 10.8% and 46.1% lower latency than TOCOM-TEM and JPEG, respectively. The encoding latency results of our PIB in different edge devices are presented in Table III. It can be observed that the feature map generation phase dominates the overall encoding latency, while the entropy coding phase contributes a negligible amount of time. Furthermore, edge devices with higher computating capacity exhibit significantly lower encoding latency.
@@ VI Conclusion
In this paper, we have proposed the Prioritized Information Bottleneck (PIB) framework as a robust solution for collaborative edge video analytics. Our contributions are two-fold. First, we have developed a prioritized inference mechanism to intelligently determine the importance of different cameras’ FOVs, effectively addressing the constraints imposed by channel capacity and data redundancy. Second, the PIB framework showcases its effectiveness by notably decreasing communication overhead and improving tracking accuracy without requiring video reconstruction at the edge server. Extensive real-world experiments show that: PIB not only surpasses the performance of conventional methods like TOCOM-TEM, JPEG, H.264, H.265, and AV1 with a marked improvement of up to 17.8% in MODA but also achieves a considerable reduction in communication costs by 82.65%, while retaining low latency and high-quality multi-view sensory data processing under less favorable channel conditions.